Hi, Can you please me solve this problem thank you very much.
The question is:
A building hold 80 apartments. Une bedroom apartment cost $850/month. A two bedroom apartment cost 1050/month. The location itself cost 78 200/month. How many one bedrooms and two bedrooms are in this building?
Thank you again.
-kim

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

X+Y=80
X=80-Y
850*X+1,050Y=78,200
NOW SUBSTITUTE
850(80-Y)+1,050Y=78,200
68,000-850Y+1,050Y=78,200
200Y=78,200-68,000
200Y=10,200
Y=10,200/200
Y=51
X+51=80
X=80-51
X=29

Answer by stanbon(75887)

A building hold 80 apartments. Une bedroom apartment cost $850/month. A two bedroom apartment cost 1050/month. The location itself cost 78200/month. How many one bedrooms and two bedrooms are in this building?
-------------------
Let number of one bedroom be "x"; cost of these is 850x dollars.
Number of two bedroom is "80-x; cost is 1050(80-x)=84000-1050x dollars
------------
EQUATION:
cost + cost = 78200
850x + 84000-1050x = 78200
-200x=-5800
x=29 (number of one-bedroom apartments)
80-x=51 (number of two-bedroom apartments)
Cheers,
Stan H.